What volume of a 2.00 M HCl solution is required to prepare 200 ml of a 0.500M solution?
1
Expert's answer
2015-11-25T09:25:16-0500
Solution:
Number of Moles of HCl in solution is:
M(HCl) = V(volume)*C(concentration) = 0.2*0.5 = 0.1 (mole);
V(HCl) = 0.1 / 2 = 0.05 (liter) = 50 ml;

Answer:
50 ml